a 720 ml container hold a mixture of red, yellow and white paint at a 5: 3: 12 ratio. How much more red paint than yellow is present in the mixture.
a. 20 ml
b. 72 ml
c. 108 ml
d. 180 ml
e. 200 ml

Answer by rfer(16322)

b) 72 ml
720/5+3+12=36
36*5=180
36*3=108
180-108=72 ml